OpenStudy (anonymous):

The ratio of the surface area of two spheres is 3:2. The volume of the larger sphere is 2,916 in3. What is the volume of the smaller sphere? 5,357 in3 1,587 in3 864 in3 1,296 in3

OpenStudy (jdoe0001):

\(\large { \cfrac{volume_{large}}{volume_{small}}=\cfrac{(side_{large})^3}{(side_{small})^3}\implies \cfrac{volume_{large}}{volume_{small}}=\left(\cfrac{(side_{large})}{(side_{small})}\right)^{\color{brown}{ 3}} \\ \quad \\ \sqrt[{\color{brown}{ 3}}]{\cfrac{volume_{large}}{volume_{small}}}=\cfrac{(side_{large})}{(side_{small})}\qquad thus \\ \quad \\ \cfrac{2,916}{x}=\left(\cfrac{3}{2}\right)^3\implies \sqrt[3]{\cfrac{2,916}{x}}=\cfrac{3}{2} }\) solve for "x"